# The Compton-Getting origin of the large-scale anisotropy of Galactic cosmic rays

### v2

@@ Abstract
Recent studies suggest that the anisotropy in cosmic-ray arrival directions can provide insight into local acceleration sites and propagation conditions. We developed a unified framework to interpret both the observed energy spectra and the large-scale anisotropy. In this work, we explore the influence of the Sun’s motion relative to the local plasma frame—the Compton-Getting (CG) effect—on the anisotropy. We find that incorporating the CG effect could reduce the dipole amplitude and slightly shift the phase away from the direction of the local regular magnetic field at tens of TeV. At lower energies, where the anisotropy from the cosmic-ray density gradient is weak, the Sun’s relative motion becomes more prominent. Below $\sim 200$ GeV, the dipole amplitude increases again, approaching the value expected from the CG effect. Additionally, a phase flip is observed at a few hundred GeV, aligning with the CG direction. Future anisotropy measurements from $100$ GeV to TeV energies could serve as a critical test of this effect.
@@ 1 Introduction
The origin and propagation of Galactic cosmic rays (CRs) below PeV energies remain unresolved questions in astrophysics. Observational data on CRs include their energy spectra, composition, and arrival directions. While earlier studies primarily focused on energy spectra and composition (Strong et al., 2007), cosmic ray (CR) anisotropy has attracted growing attention in recent years. Due to frequent scattering by interstellar magnetohydrodynamic (MHD) turbulence, the arrival directions of CRs are nearly isotropic (Jokipii, 1966; Schlickeiser, 2002; Yan, 2022). However, numerous measurements have revealed a nonuniform distribution of arrival directions, known as anisotropy (Amenomori et al., 2006; Guillian et al., 2007; Abdo et al., 2008, 2009; Abbasi et al., 2010; Amenomori et al., 2010; Abbasi et al., 2011, 2012; Aartsen et al., 2013; Bartoli et al., 2013; Abeysekara et al., 2014; Bartoli et al., 2015; Aartsen et al., 2016; Amenomori et al., 2017; López-Barquero et al., 2017; Abeysekara et al., 2019). The intensity of anisotropy varies between $\sim 10^{-4}$ and $\sim 10^{-3}$ in the TeV–PeV range and exhibits energy-dependent angular distributions.
@@
Despite the fact that the CR trajectories are highly tangled by the interstellar magnetic field, abundant studies have shown that anisotropy can still retain the information of CR sources. Within the framework of the traditional diffusion model, the uneven distribution of overall CR sources predicts a large-scale anisotropy, with the direction aligning with the Galactic center immutably(Blasi & Amato, 2012; Liu et al., 2017). If a local source contributes significantly to the CR flux, anisotropy may indicate its presence (Manconi et al., 2017; Luo et al., 2022). In recent work, we developed a unified scenario that incorporatesates a local source within a spatially dependent propagation model to explain both the energy specand the and the dipole anisotropy (Liu et al., 2019; Qiao et al., 2019, 2023). This model suggests that the excess of nuclei between 200 GeV and $\sim 20$ TeV, along with the energy-dependent anisotropy below 100 TeV, implies the presence of nearby CR sources. The supernova remnant (SNR) at the birthplace of the Geminga pulsar could be a plausible candidate for such a local cosmic ray accelerator(Zhao et al., 2022), with similar conclusions supported by other recent studies (Zhang et al., 2022b; Nie et al., 2023).
@@
Anisotropy also provides indirect insight into the local interstellar environment. Observations of energetic neutral atom emissions by the IBEX mission reveal a local regular magnetic field (LRMF) within 0.1 pc of the solar system, oriented at $l\simeq 210.5^{\circ}$, $b\simeq-57.1^{\circ}$ (McComas et al., 2009). This direction closely matches the observed dipole phase below 100 TeV. Studies suggest that anisotropic diffusion guided by the LRMF directs CRs into the heliosphere (Schwadron et al., 2014; Ahlers, 2016; Liu et al., 2020). Furthermore, the energy dependence of the dipole phase implies a transition from anisotropic to isotropic diffusion around 100 TeV in the local interstellar environment(Li et al., 2024). Moreover, quadrupole anisotropies may be induced by nonuniform convection effects (Zhang et al., 2022a), while smaller-scale anisotropies are likely attributed to the local turbulence on the scale of a few mean free paths, or heliospheric influences. For a comprehensive review, one can refer to Ahlers & Mertsch (2017).
@@
Notably, the non-uniform distribution of the arrival directions of CRs was predicted as early as the 1930s by Compton & Getting (1935). This is caused by the relative motion of the solar system through the local plasma frame. As the Sun has an average circular speed $v\simeq 220$ km/s around the Galactic center, it induces a dipole anisotropy, which is similar to the Doppler effect. The amplitude of this anisotropy is related to the power index of the CR spectrum and is expected to be $2.3\times 10^{-3}$ with its peak at $\alpha\simeq 315^{\circ}$, $\delta\simeq 0^{\circ}$. This is well known as the Compton-Getting (CG) effect. However, the expected dipole anisotropy is inconsistent with the measurement at $\sim 300$ TeV (Amenomori et al., 2006). Thereupon, it is inferred that the Galactic CRs co-rotate with the local Galactic magnetic field environment. Nevertheless, another CG effect, which occurs in local solar time, has been detected. It originates from the evolution of the Earth around the Sun (Cutler & Groom, 1986; Amenomori et al., 2004). Since the revolution speed of the Earth is $\sim 30~{}$km/s, the amplitude is ${}4.5\times 10^{-4}$ and points to the local $6$ hour a.m. For the ultra-high energy cosmic rays, most of them are regarded as the extragalactic origin. It has been argued that the motion of the Sun relative to the cosmic microwave background (CMB) frame could induce a dipole anisotropy with intensity $\sim 0.6\%$, similar to that seen in the CMB (Kachelrieß & Serpico, 2006).
@@
In this work, we further investigate the impact of the CG effect on CR anisotropy. Although CRs seem to corotate with the average motion of the stars in the whole Galaxy, the motion of the Sun relative to the local plasma frame is considered. Since the plasma rest frame is not precisely defined, we explore two scenarios: one is the so-called local standard of rest (LSR), which corresponds to the Sun’s motion towards the solar apex, and the other is the relative motion through the local interstellar medium. We find that in both cases, below 10 TeV, the dipole phase no longer points towards the direction of the local regular magnetic field at the lower energy as the anisotropy from the CR gradient gradually decreases. Our predictions can be tested by future anisotropy measurements at sub-TeV energies(Cao et al., 2019; Pan et al., 2024).
@@ 2.1 Spatially-dependent propagation
In this work, we adopt the spatial-dependent propagation (SDP) model. Over the past decade, this model has been extensively applied to cosmic-ray (CR) propagation to interpret a wide range of observational data (Tomassetti, 2012, 2015; Feng et al., 2016; Guo et al., 2016; Liu et al., 2018; Guo & Yuan, 2018; Liu et al., 2019; Qiao et al., 2019; Tian et al., 2020; Yuan et al., 2020; Zhao et al., 2022). Recent observations of TeV halos surrounding pulsars (Abeysekara et al., 2017; Aharonian et al., 2021; H. E. S. S. Collaboration et al., 2023; Albert et al., 2024), along with evidence of anisotropic diffusion in strongly magnetized media (i.e., with $M_{A}\simeq\delta B/B_{0}\leq 1$) (Yan & Lazarian, 2008; Cerri et al., 2017; Liu et al., 2020; Maiti et al., 2022), strongly support the necessity of incorporating spatial dependence into the diffusion process (Yan & Lazarian, 2002, 2004; Evoli & Yan, 2014).
In our SDP model, the entire diffusive halo is divided into two regions. The region near the Galactic disk and its immediate surroundings is termed the inner halo (IH), where turbulence is significantly influenced by supernova activity. Consequently, areas with higher source density exhibit stronger turbulence and slower diffusion, whereas regions with fewer sources have more moderate turbulence and faster diffusion. As a result, the diffusion coefficient in the IH depends on the radial distribution of CR sources. In contrast, the outer halo (OH), located farther from the Galactic disk, is less affected by stellar activity, and the diffusion coefficient there depends only on particle rigidity.
@@
where ${\cal R}_{0}$ is the reference rigidity fixed at 4 GV, $\beta_{0}$ is the particle velocity normalized to the speed of light, $D_{0}$ and $\delta_{0}$ are constants representing the diffusion coefficient and its rigidity dependence in the OH, and $\eta$ is a phenomenological parameter introduced to fit low-energy data. The spatial function $F(r,z)$ is given by:
$F(r,z)=\begin{cases}g(r,z)+\left[1-g(r,z)\right]\left(\dfrac{z}{\xi z_{h}}\right)^{n},&|z|\leqslant\xi z_{h}\\ 1~{},&|z|>\xi z_{h}\end{cases}.$ (2)
where $g(r,z)=N_{m}/[1+f(r,z)]$, with $f(r,z)$ representing the spatial distribution of CR sources. The CR sources are assumed to be axisymmetrically distributed as $f(r,z)\propto(r/r_{\odot})^{\alpha}\exp[-\beta(r-r_{\odot})/r_{\odot}]\exp(-|z|/z_{s})$, where $r_{\odot}=8.5$ kpc and $z_{s}=0.2$ kpc. $\alpha$ and $\beta$ are taken as $1.69$ and $3.33$ (Case & Bhattacharya, 1996).
@@
The diffusion-reacceleration (DR) model is adopted in the propagation equation and the numerical package, DRAGON, is used to compute the background CR distribution (Evoli et al., 2008).
@@ 2.2 Local source
In our previous work (Liu et al., 2019; Qiao et al., 2019; Zhao et al., 2022), we attributed the spectral hardening above 200 GV and the dipole anisotropy below 100 TeV to the Geminga supernova remnant (SNR), which resulted from the explosion of the Geminga progenitor. The characteristic age of the SNR, inferred from the Geminga pulsar’s spin-down luminosity, is about $\tau=3.4\times 10^{5}$ years (Manchester et al., 2005). Its inferred position is $l=194.3^{\circ},b=-13^{\circ}$, and it lies approximately 330 pc from the Solar System (Faherty et al., 2007). The CR flux from the Geminga SNR can be computed by solving the time-dependent diffusion equation using the Green’s function method, assuming a boundary at infinity (Liu et al., 2017, 2019). As for the instantaneous and point-like injection, the CR density of local source can be calculated as
@@
where $\sigma({\cal R},t)=\sqrt{2D_{xx}({\cal R})t}$ is the effective diffusion length within time $t$. The injection spectrum $q_{\rm inj}({\cal R})$ is parameterized as a power-law function of rigidity with an exponential cutoff, i.e. $q_{0}{\cal R}^{-\alpha}\exp(-{\cal R}/{\cal R}^{\prime}_{\rm c})$.
@@ 2.3 Anisotropic diffusion
The IBEX experiment has detected a local regular magnetic field (LRMF) with a spatial scale of $\sim 0.1$ pc surrounding the Solar System (McComas et al., 2009; Funsten et al., 2013), oriented along $l\simeq 210.5^{\circ},b\simeq-57.1^{\circ}$. Similar results have been reported via polarization measurements of nearby stars (Frisch et al., 2015). This direction coincides with the observed dipole phase below 100 TeV (Ahlers, 2016). Since the spatial scale of LRMF is much smaller than the average propagation length inferred from the boron-to-carbon ratio, the LRMF does not have a remarkable impact on the energy spectra. However, in the local interstellar medium, CRs propagate anisotropically due to the influence of LRMF, which modifies their arrival direction and dipole anisotropy correspondingly. In this scenario, the scalar diffusion coefficient $D_{xx}$ is replaced by a diffusion tensor $D_{ij}$, expressed as:
$D_{ij}=D_{\perp}\delta_{ij}+(D_{\parallel}-D_{\perp})b_{i}b_{j}~{},\quad b_{i}=\dfrac{B_{i}}{|\boldsymbol{B}|}$ (5)
where $b_{i}$ is the $i$-th component of the LRMF unit vector (Giacalone & Jokipii, 1999; Cerri et al., 2017). The diffusion coefficients parallel and perpendicular to the LRMF are given by:
@@
$D_{\perp}$ $=\,D_{0\perp}\beta_{0}^{\eta}\left(\frac{\cal R}{{\cal R}_{0}}\right)^{\delta_{\perp}}=\varepsilon D_{0\parallel}\beta_{0}^{\eta}\left(\frac{\cal R}{{\cal R}_{0}}\right)^{\delta_{\perp}}~{}.$ (7)
where $D_{\parallel}=D_{xx}$, i.e., $D_{0\parallel}=D_{0}F(r,z)$ and $\delta_{\parallel}=\delta_{0}F(r,z)$. $\varepsilon=D_{0\perp}/D_{0\parallel}$ is the ratio between perpendicular and parallel diffusion coefficient at the reference rigidity ${\cal R}_{0}$. The energy dependence of the ratio between perpendicular and parallel diffusion coefficients may be attributed to the difference in the path volumes that particles of different energy go through. The higher energy particles propagating through the outer or inner halos before reaching the solar system are like to traverse a larger path volume, which includes more sources, and thus experience stronger perturbations. In contrast, low-energy particles travel through a smaller path volume and naturally undergo weaker disturbances. This could lead to an energy dependence in this ratio. Taking local anisotropic diffusion into account, the dipole anisotropy becomes:
$\boldsymbol{\delta}=\dfrac{3\boldsymbol{D}\cdot\nabla\psi}{v\psi}=\dfrac{3}{v\psi}D_{ij}\dfrac{\partial\psi}{\partial x_{j}}.$ (8)

### v3

@@ Abstract
Recent studies suggest that the anisotropy in cosmic-ray arrival directions can provide insight into local acceleration sites and propagation conditions. We developed a unified framework to interpret both the observed energy spectra and the large-scale anisotropy. In this work, we explore the influence of the Sun’s motion relative to the local plasma frame—the Compton-Getting (CG) effect—on the anisotropy. We find that incorporating the CG effect could slightly reduce the dipole amplitude and shift the phase away from the direction of the local regular magnetic field at tens of TeV. At lower energies, where the anisotropy from the cosmic-ray density gradient is weak, the Sun’s relative motion becomes more prominent. Below $\sim 200$ GeV, the dipole amplitude increases again, approaching the value expected from the CG effect. Additionally, a phase flip is observed at a few hundred GeV, aligning with the CG direction. Future anisotropy measurements from $100$ GeV to TeV energies could serve as a critical test of this effect.
@@ 1 Introduction
The origin and propagation of Galactic cosmic rays (CRs) below PeV energies remain unresolved questions in astrophysics. Observational data on CRs include their energy spectra, composition, and arrival directions. While earlier studies primarily focused on energy spectra and composition (2007ARNPS..57..285S), cosmic-ray (CR) anisotropy has attracted growing attention in recent years. Due to frequent scattering by interstellar magnetohydrodynamic (MHD) turbulence, the arrival directions of CRs are nearly isotropic (1966ApJ...146..480J; 2002cra..book.....S; Yan2022icrc). However, numerous measurements have revealed a nonuniform distribution of arrival directions, known as anisotropy (2006Sci...314..439A; 2007PhRvD..75f2003G; 2008PhRvL.101v1101A; 2009ApJ...698.2121A; 2010ApJ...718L.194A; 2010ApJ...711..119A; 2011ApJ...740...16A; 2012ApJ...746...33A; 2013ApJ...765...55A; 2013PhRvD..88h2001B; 2014ApJ...796..108A; 2015ApJ...809...90B; 2016ApJ...826..220A; 2017ApJ...836..153A; 2017ApJ...842...54L; 2019ApJ...871...96A). The intensity of anisotropy varies between $\sim 10^{-4}$ and $\sim 10^{-3}$ in the TeV–PeV range and exhibits energy-dependent angular distributions.
@@
Despite the fact that the CR trajectories are highly tangled by the interstellar magnetic field, abundant studies have shown that anisotropy can still retain some information of CR sources. Within the framework of the traditional diffusion model, the uneven distribution of overall CR sources predicts a large-scale anisotropy, with the direction aligning with the Galactic center immutably(2012JCAP...01..011B; 2017PhRvD..96b3006L). If a local source contributes significantly to the CR flux, anisotropy may indicate its presence (2017JCAP...01..006M; 2022ApJ...930...82L). In recent works, we developed a unified scenario that incorporates a local source within a spatially dependent propagation model to explain both the energy spectrum and the dipole anisotropy (2019JCAP...10..010L; 2019JCAP...12..007Q; 2023ApJ...942...13Q). This model suggests that the excess of nuclei between 200 GeV and $\sim 20$ TeV, along with the energy-dependent anisotropy below 100 TeV, implies the presence of nearby CR sources. The supernova remnant (SNR) at the birthplace of the Geminga pulsar could be a plausible candidate for such a local cosmic ray accelerator(2022ApJ...926...41Z), with similar conclusions supported by other recent studies (2022MNRAS.511.6218Z; 2023ApJ...952..100N).
@@
The anisotropy also provides indirect insight into the local interstellar environment. Observations of energetic neutral atom emissions by the IBEX mission reveal a local regular magnetic field (LRMF) within 0.1 pc of the solar system, oriented at $l\simeq 210.5^{\circ}$, $b\simeq-57.1^{\circ}$ (2009Sci...326..959M). This direction closely matches the observed dipole phase below 100 TeV. Some studies suggest that anisotropic diffusion guided by the LRMF directs CRs into the heliosphere (2014Sci...343..988S; 2016PhRvL.117o1103A; 2020ApJ...892....6L). Furthermore, the energy-dependent of the dipole phase implies a transition from anisotropic to isotropic diffusion around 100 TeV in the local interstellar environment(2024ApJ...962...43L). Moreover, quadrupole anisotropies may be induced by nonuniform convection effects (2022ApJ...938..106Z), while smaller-scale anisotropies are likely attributed to the local turbulence on the scale of a few mean free paths, or heliospheric influences. For a comprehensive review, one can refer to 2017PrPNP..94..184A.
@@
Notably, the non-uniform distribution of the arrival directions of CRs was predicted as early as the 1930s by 1935PhRv...47..817C. This is caused by the relative motion of the solar system through the local plasma frame. As the Sun has an average circular speed $v\simeq 220$ km/s around the Galactic center, it induces a dipole anisotropy, which is similar to the Doppler effect. The amplitude of this anisotropy is related to the power index of the CR spectrum and is expected to be $2.3\times 10^{-3}$ with its peak at $\alpha\simeq 315^{\circ}$, $\delta\simeq 0^{\circ}$. This is well known as the Compton-Getting (CG) effect, denoted as $\rm CG_{GAL}$. However, the expected dipole anisotropy is inconsistent with the measurement at $\sim 300$ TeV (2006Sci...314..439A). Thereupon, it is inferred that the Galactic CRs co-rotate with the local Galactic magnetic field environment. Nevertheless, another CG effect ($\rm CG_{SUN}$), which occurs in local solar time, has been detected. It originates from the evolution of the Earth around the Sun (1986Natur.322..434C; 2004PhRvL..93f1101A). Since the revolution speed of the Earth is $\sim 30$~km/s, the amplitude is $4.5\times 10^{-4}$ and points to the local $6$ hour a.m. For the ultra-high energy cosmic rays, most of them are regarded of extragalactic origin. It has been argued that the motion of the Sun relative to the cosmic microwave background (CMB) frame could induce a dipole anisotropy with intensity $\sim 0.6\%$, similar to that seen in the CMB (2006PhLB..640..225K).
@@
In this work, we further investigate the impact of the CG effect, including $\rm CG_{LSR}$ and $\rm CG_{ISM}$, on CR anisotropy. Although CRs seem to corotate with the average motion of the stars in the whole Galaxy, the motion of the Sun relative to the local plasma frame is considered. Since the plasma rest frame is not precisely defined, we explore two scenarios: one is the so-called local standard of rest (LSR), which corresponds to the Sun’s motion towards the solar apex, and the other is the relative motion through the local interstellar medium(ISM). We find that in both cases, below 10 TeV, the dipole phase no longer points towards the direction of the local regular magnetic field at the lower energy as the anisotropy from the CR gradient gradually decreases. Our predictions can be tested by future anisotropy measurements at sub-TeV energies(2019arXiv190502773C; 2024NuScT..35..149P).
@@ 2.1 Spatially-dependent propagation
In this study, we employ the spatially dependent propagation (SDP) model to describe Galactic cosmic-ray transport. This class of models has been developed and applied extensively over the past decades, providing a unified framework for interpreting diverse observational results, such as the hardening of primary and secondary spectra, diffuse $\gamma$-ray emission, and anisotropy at large scales. (2012ApJ...752L..13T; 2015PhRvD..92h1301T; 2016PhRvD..94l3007F; 2016ApJ...819...54G; 2018ApJ...869..176L; 2018PhRvD..97f3008G; 2019JCAP...10..010L; 2019JCAP...12..007Q; 2020ChPhC..44h5102T; 2020FrPhy..1624501Y; 2022ApJ...926...41Z). Recent observations of TeV halos surrounding pulsars (2017Sci...358..911A; 2021PhRvL.126x1103A; 2023A&A...673A.148H; 2024ApJ...974..246A), along with evidence of anisotropic diffusion in strongly magnetized turbulence (i.e., with $M_{A}\simeq\delta B/B_{0}\leq 1$) (2008ApJ...673..942Y; 2017JCAP...10..019C; 2020ApJ...892....6L; 2022ApJ...926...94M), also strongly support the necessity of incorporating spatial dependence into the diffusion process, as demonstrated earlier from the perspective of inhomogeneous turbulence driving and damping (2002PhRvL..89B1102Y; 2004ApJ...614..757Y; 2014ApJ...782...36E; 2020NatAs...4.1001Z; 2020PhRvX..10c1021M; Hou2025; Zhao2025).
In the SDP picture, the Galactic halo is conceptually divided into two transport layers. The inner halo (IH), extending around the disk, is strongly affected by supernova-driven turbulence, which suppresses diffusion in regions of high source density. Conversely, in regions with relatively few sources, turbulence levels are lower and diffusion proceeds more efficiently. As a result, the diffusion coefficient in the IH depends on the radial distribution of CR sources. In contrast, the outer halo (OH), located farther from the Galactic disk, is less affected by stellar activity, and the diffusion coefficient there depends only on particle rigidity.
@@
where ${\cal R}_{0}$ is the reference rigidity fixed at 4 GV, $\beta_{0}$ is the particle velocity normalized to the speed of light, $D_{0}$ and $\delta_{0}$ are constants representing the diffusion coefficient and its rigidity-dependent in the OH, and $\eta$ is a phenomenological parameter introduced to fit low-energy data. The spatial function $F(r,z)$ is given by:
$F(r,z)=\begin{cases}g(r,z)+\left[1-g(r,z)\right]\left(\dfrac{z}{\xi z_{h}}\right)^{n},&|z|\leqslant\xi z_{h}\\ 1~,&|z|>\xi z_{h}\end{cases},$ (2)
where $g(r,z)=N_{m}/[1+f(r,z)]$, with $f(r,z)$ representing the spatial distribution of CR sources. The CR sources are assumed to be axisymmetrically distributed as $f(r,z)\propto(r/r_{\odot})^{\alpha}\exp[-\beta(r-r_{\odot})/r_{\odot}]\exp(-|z|/z_{s})$, where $r_{\odot}=8.5$ kpc and $z_{s}=0.2$ kpc. $\alpha$ and $\beta$ are taken as $1.69$ and $3.33$ (1996A&AS..120C.437C).
@@
The diffusion-reacceleration (DR) model is adopted in the propagation equation and the numerical package, DRAGON, is used to compute the background CR distribution (1475-7516-2008-10-018).
@@ 2.2 Local source
Building on earlier analyses (2019JCAP...10..010L; 2019JCAP...12..007Q; 2022ApJ...926...41Z), we consider the Geminga supernova remnant (SNR) as the likely origin of the spectral hardening above a few hundred GV and of the dipole anisotropy observed below $\sim$100 TeV. The age of the remnant, estimated from the spin-down evolution of the Geminga pulsar, is about $\tau=3.4\times 10^{5}$ years (2005AJ....129.1993M). Its location is determined to be $l=194.3^{\circ},b=-13^{\circ}$, at a distance of approximately 330 pc from the Sun (2007Ap&SS.308..225F). The contribution of this nearby source is modeled by solving the time-dependent diffusion equation using the Green’s function formalism, with free-escape conditions at large distances (2017PhRvD..96b3006L; 2019JCAP...10..010L). For the instantaneous and point-like injection, the CR density of local source can be calculated as
@@
where $\sigma({\cal R},t)=\sqrt{2D_{xx}({\cal R})t}$ is the effective diffusion length within time $t$. The injection spectrum $q_{\rm inj}({\cal R})$ is parameterized as a power-law function of rigidity with an exponential cutoff, i.e. $q_{0}{\cal R}^{-\alpha}\exp(-{\cal R}/{\cal R}^{\prime}_{\rm c})$. Note that the Geminga nearby source, which located at a specific position in the inner halo, is very close to the Solar System. Hence, its diffusion coefficient is taken as the local value at the Solar System given in Eq. (1) here.
Based on the injection spectrum parameters listed in Tab. 3 and the formula $W_{p,He}=\int q_{inj}(E)EdE$, we obtain total injection energies of about $1.5\times 10^{50}$ erg for protons and $6.8\times 10^{49}$ erg for helium from nearby source, corresponding to roughly 5-15% of a canonical supernova explosion energy of $10^{51}$ erg, consistent with standard cosmic-ray acceleration efficiencies. For the background component, the estimated energy injection that includes all Galactic cosmic-ray sources except Geminga is approximately $6.5\times 10^{50}$ erg, which is physically comparable to the injection energy from the nearby source. Since the background and nearby source components are computed using the numerical software(DRAGON) and the analytical Green’s function method, respectively, different units are therefore required to ensure dimensional consistency here.
Why do we not include other nearby sources such as Monogem, Cygnus Loop, and Vela? Because their contributions to the observed proton spectrum are strongly suppressed by their age-distance combinations. To illustrate this point, we separately computed the proton spectra from these nearby sources. Their corresponding parameters are summarized in Tab. 1. For Monogem, the injection energy is scaled according to its X-ray–inferred explosion energy of $\sim$ $1.9\times 10^{50}$ erg (1996ApJ...463..224P), which is about one order of magnitude lower than that of Geminga. For Vela and Cygnus Loop, the injection energies are assumed to be comparable to that of Geminga to facilitate a consistent comparison. These parameter selections are guided by observational and physical constraints, ensuring that the model predictions are well grounded rather than arbitrary. Under the same transport setup, our calculations show that the predicted proton fluxes from these nearby sources at Earth remain sub-dominant compared to that of Geminga. For relatively young sources, low-energy particles have not yet had sufficient time to diffuse to the Solar system, while for the more distant ones, the local flux is significantly suppressed. A systematic study by 2022ApJ...930...82L also demonstrated that, under these considerations, Geminga stands out as the most plausible nearby contributor, while the other candidates are either negligible or inconsistent with current spectral and anisotropy data.
Tab. 1: Characteristic and injection parameters for other three nearby SNRs.
  SNR Name   Distance (pc)   Age (yr)      $q_{0}^{p}~[\rm GeV^{-1}]$      $\alpha$      ${\cal R}^{\prime}_{c}$ [TV]
  Monogem   288   $8.6\times 10^{4}$   $3.2\times 10^{51}$   2.10   30
  Vela   300   $1.1\times 10^{4}$   $3.2\times 10^{52}$   2.10   30
  Cygnus Loop   720   $2.0\times 10^{4}$   $3.2\times 10^{52}$   2.10   30
@@ 2.3 Anisotropic diffusion
Observations with IBEX reveal a local regular magnetic field (LRMF) in the solar neighborhood with a characteristic spatial scale of order $\sim 0.1$ pc and orientation $l\simeq 210.5^{\circ},b\simeq-57.1^{\circ}$ (2009Sci...326..959M; 2013ApJ...776...30F). Starlight polarization measurements independently support a consistent field direction (2015ApJ...814..112F), which is broadly aligned with the measured dipole phase below $\sim$100 TeV (2016PhRvL.117o1103A). Because the spatial scale of the LRMF ($\sim$0.1 pc) is much shorter than the typical propagation scale of Galactic CRs inferred from the B/C ratio, its influence on the energy spectra is negligible. In other words, the large-scale spectral features are governed primarily by transport over kiloparsec distances and are insensitive to magnetic-field structures on sub-parsec scales. Nevertheless, such a field can have a pronounced effect on the angular distribution of arriving particles, since the local anisotropy is shaped by the most recent segment of the propagation history, where the guiding effect of the LRMF can redirect particle trajectories without altering their overall energy distribution. In this scenario, the scalar diffusion coefficient $D_{xx}$ is replaced by a diffusion tensor $D_{ij}$, expressed as:
$D_{ij}=D_{\perp}\delta_{ij}+(D_{\parallel}-D_{\perp})b_{i}b_{j}~,\quad b_{i}=\dfrac{B_{i}}{|\boldsymbol{B}|}~,$ (5)
where $b_{i}$ is the $i$-th component of the LRMF unit vector (1999ApJ...520..204G; 2017JCAP...10..019C). The diffusion coefficients parallel and perpendicular to the LRMF are given by:
@@
$D_{\perp}$ $=\,D_{0\perp}\beta_{0}^{\eta}\left(\frac{\cal R}{{\cal R}_{0}}\right)^{\delta_{\perp}}=\varepsilon D_{0\parallel}\beta_{0}^{\eta}\left(\frac{\cal R}{{\cal R}_{0}}\right)^{\delta_{\perp}}~,$ (7)
where $D_{\parallel}=D_{xx}$, i.e., $D_{0\parallel}=D_{0}F(r,z)$ and $\delta_{\parallel}=\delta_{0}F(r,z)$. $\varepsilon=D_{0\perp}/D_{0\parallel}$ is the ratio between perpendicular and parallel diffusion coefficient at the reference rigidity ${\cal R}_{0}$. The energy-dependent of the ratio between perpendicular and parallel diffusion coefficients may be attributed to the difference in the path volumes that particles of different energy go through. The higher energy particles propagating through the outer or inner halos before reaching the solar system are like to traverse a larger path volume, which includes more sources, and thus experience stronger perturbations. In contrast, low-energy particles travel through a smaller path volume and naturally undergo weaker disturbances. This could lead to an energy-dependent in this ratio. Taking local anisotropic diffusion into account, the dipole anisotropy becomes:
$\boldsymbol{\Delta_{grad}}=\dfrac{3\boldsymbol{D}\cdot\nabla\psi}{v\psi}=\dfrac{3}{v\psi}D_{ij}\dfrac{\partial\psi}{\partial x_{j}}.$ (8)